what is the length of the diagonal of a cube with a surface area of 24units squared?

Answer by mananth(16946)

the surface area of cube = 24
Let side be x
area of one face of cube = x^2
Total surface area = area of 6 faces
=6x^2
6x^2=24
x^2=4
x=2
..
Use pythagoras theorem to find diagonal D
2^2 +2^2 = D62
8=D^2
d= sqrt 8